# Waterfall: Framework for Robust and Scalable Text Watermarking

### v1

@@ 7 Limitations
As Waterfall relies on the adjustment of the original text to add watermarks, it may not applicable to all types of text. For example, Waterfall faces limitations in its application to works where their IP values lie in their style or format (e.g., poems), unless additional methods are applied that cause LLMs to largely preserve such styles while paraphrasing these works, such as optimizing for better paraphrasing prompts to be used with more capable LLMs, or iteratively refining the text through multiple rounds of watermarking.
@@ Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD/2023-01-039J). This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. Xinyuan Niu is supported by the Centre for Frontier AI Research of Agency for Science, Technology and Research (ASTAR). Jiangwei Chen is supported by the Institute for Infocomm Research of Agency for Science, Technology and Research (ASTAR). We acknowledge CSC (Finland) for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, and hosted by CSC (Finland) and the LUMI consortium. The access was made possible via collaboration between NSCC (Singapore) and CSC (Finland).
@@ E.5 Fidelity
We provide some examples of text watermarked by the Waterfall, M-bit and P-nlw. LABEL:tab:c4_samples shows a few samples from the c4 dataset with watermarked text of varying STS scores. M-bit has the highest STS across these samples listed, due to its algorithm only changing very few words within the text, resulting in lower scalability as described in Section 4.1. Despite the high STS score, it can be visually seen that text watermarked with M-bit and P-nlw introduces linguistic and grammatical errors to the text, which are not measured by the STS score.
@@
E.6 Verifiability fidelity trade-off
@@ E.7 Scalability
E.7.1 Scalability when verifying with different IDs
@@
E.7.2 Scalability when watermarking different IDs
@@
E.7.3 Discussion on scalability in practice
@@ F.3 $\mathbb{A}_{3}$
We consider the worst-case scenario of robust verifiability under $\mathbb{A}_{3}$ for two traditional approaches P-nlw and M-bit. Because these two methods are based on embedding binary keys in the watermarking stage, we try to apply $\mathbb{A}_{3}$ with the complement of the binary watermark key that was extracted as part of the verification process (replacing bit 0 with bit 1 and vice versa), to illustrate the worst-case scenario. We conduct this experiment with setting as 4.1. The results are illustrated in Table 7 and Figure 16. Do note that attacks could engineer their attacks by performing overlap watermarking with a mixture of watermark bits, random bits and complement bits, to target any AUROC value between the pre-attack and overlap complement AUROC.
@@ G.1 Code watermarking experiment settings
In the main paper, we report the result of code watermarking on the MBJSP dataset (Athiwaratkun et al., 2023) with the data ownership problem setting. This is a JavaScript dataset including around 800 crowd-sourced JavaScript programming problems. To show the ability of Waterfall on watermarking other programming languages, we also perform data ownership watermarking on Python datasets, which can be found in Appendix G.5.
@@ Appendix I Adapting model watermarking schemes into Waterfall framework
To the best of our knowledge, the only work related to this topic we have found so far is the unpublished work (Yang et al., 2023) which applies text watermarking methods to the specific use case of text generated by black-box language models and is therefore essentially a text watermarking paper. The text watermarking method of Yang et al. (2023) is similar to the M-BIT benchmark (Yoo et al., 2023) that we considered in the main paper, and essentially encodes watermarks by first identifying words to replace (based on linguistic rules), then finds synonyms for them which are used to represent bits of the watermarking signal. Although the two methods differ in the way of selecting which word to perform watermarking (sentence/word embedding similarity for Yang et al. (2023) and a 2nd BERT model for M-BIT), given their similar characteristics, both methods ultimately still suffer from robust verifiability compared to Waterfall.
@@
We ran additional experiments to compare this Waterfall variant [Waterfall (KGW)] with our original watermarking scheme [Waterfall (Ours)]. Figure 21 demonstrate that Waterfall (Ours) has a strictly better fidelity-verifiability Pareto frontier, i.e., for any required fidelity (STS score), Waterfall (Ours) has higher verifiability than Waterfall (KGW).

### v2

@@ 7 Limitations
As Waterfall relies on the adjustment of the original text to add watermarks, it may not applicable to all types of text. For example, Waterfall faces limitations in its application to works where their IP values lie in their style or format (e.g., poems), unless additional methods are applied that cause LLMs to largely preserve such styles while paraphrasing these works, such as optimizing for better paraphrasing prompts to be used with more capable LLMs, or iteratively refining the text through multiple rounds of watermarking. We further discuss examples of such methods in Appendix L, and also demonstrate an example of using the Reflexion technique Shinn et al. (2023) to overcome similar issues for code in Section G.2.
@@ Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD/2023-01-039J). This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This research is supported by the National Research Foundation Singapore and the Singapore Ministry of Digital Development and Innovation, National AI Group under the AI Visiting Professorship Programme (award number AIVP-$2024$-$001$). Xinyuan Niu is supported by the Centre for Frontier AI Research of Agency for Science, Technology and Research (ASTAR). Jiangwei Chen is supported by the Institute for Infocomm Research of Agency for Science, Technology and Research (ASTAR). We acknowledge CSC (Finland) for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, and hosted by CSC (Finland) and the LUMI consortium. The access was made possible via collaboration between NSCC (Singapore) and CSC (Finland).
@@ E.5 Fidelity
We provide some examples of text watermarked by the Waterfall, M-bit and P-nlw. LABEL:tab:c4_samples shows a few samples from the c4 dataset with watermarked text of varying STS scores. M-bit has the highest STS across these samples listed, due to its algorithm only changing very few words within the text, resulting in lower scalability as described in Section 4.1. Despite the high STS score, it can be visually seen that text watermarked with M-bit and P-nlw introduces linguistic and grammatical errors to the text, which are not measured by the STS score. Further analysis on the text quality is included in Section E.6.
@@
E.6 Text quality
To evaluate the fluency and grammatical corectness of the text, we used OpenAI’s gpt-4o as an evaluator to judge the text quality of the watermarked text comparing Waterfall and the benchmarks M-bit, P-nlw. Inspired by Wang et al. (2023a), we prompted gpt-4o with the following prompt.
Given the text below, give a score between 0 and 100 for the fluency and grammatical correctness of the text. Be strict and only give high scores sparingly. Only output the number and do not output any other text or explanation.
We also evaluated text quality using GPTScore (Fu et al., 2024), which is a metric to evaluate the quality and fluency using the GPT3 babbage-002 backbone.
In Figure 13, we show that benchmarks M-bit and P-nlw fall well within the Pareto frontier of Waterfall when considering text quality vs. verifiability trade-off.
Refer to captionRefer to caption
Figure 13: Text quality against verifiability trade-off. Left: Fluency score (modified from Wang et al. (2023a)) and Right: GPTScore (Fu et al., 2024)
E.7 Verifiability fidelity trade-off
@@ E.8 Scalability
E.8.1 Scalability when verifying with different IDs
@@
E.8.2 Scalability when watermarking different IDs
@@
E.8.3 Discussion on scalability in practice
@@ F.3 $\mathbb{A}_{3}$
We consider the worst-case scenario of robust verifiability under $\mathbb{A}_{3}$ for two traditional approaches P-nlw and M-bit. Because these two methods are based on embedding binary keys in the watermarking stage, we try to apply $\mathbb{A}_{3}$ with the complement of the binary watermark key that was extracted as part of the verification process (replacing bit 0 with bit 1 and vice versa), to illustrate the worst-case scenario. We conduct this experiment with setting as Section 4.1. The results are illustrated in Table 7 and Figure 18. Do note that attacks could engineer their attacks by performing overlap watermarking with a mixture of watermark bits, random bits and complement bits, to target any AUROC value between the pre-attack and overlap complement AUROC.
@@ G.1 Code watermarking experiment settings
In the main paper, we report the result of code watermarking on the MBJSP dataset (Athiwaratkun et al., 2023) with the data ownership problem setting. This is a JavaScript dataset including around 800 crowd-sourced JavaScript programming problems. To show the ability of Waterfall on watermarking other programming languages, we also perform data ownership watermarking on Python datasets, which can be found in Section G.5.
@@ Appendix I Adapting model watermarking schemes into Waterfall framework
To the best of our knowledge, the only works related to this topic we have found so far are the unpublished works (Yang et al., 2023) and (Chang et al., 2024).
Yang et al. (2023) applies text watermarking methods to the specific use case of text generated by black-box language models and is therefore essentially a text watermarking paper. The text watermarking method of Yang et al. (2023) is similar to the M-BIT benchmark (Yoo et al., 2023) that we considered in the main paper, and essentially encodes watermarks by first identifying words to replace (based on linguistic rules), then finds synonyms for them which are used to represent bits of the watermarking signal. Although the two methods differ in the way of selecting which word to perform watermarking (sentence/word embedding similarity for Yang et al. (2023) and a 2nd BERT model for M-BIT), given their similar characteristics, both methods ultimately still suffer from robust verifiability compared to Waterfall.
@@
Chang et al. (2024) is a concurrent unpublished work that applies the watermark to LLM generated text by prompting another LLM to insert a list of words into the text. The list of words are selected based on a pseudo-random mapping of text embedding to words and top-k words semantically similar to the text. Verification is performed by counting the proportion of words in the list that are present in the text. Robustness against paraphrasing is achieved by also including words from the text that are sufficiently semantically similar to the words in the list.
@@
We ran additional experiments to compare this Waterfall variant [Waterfall (KGW)] with our original watermarking scheme [Waterfall (Ours)] in Appendix I. Figure 23 demonstrate that Waterfall (Ours) has a strictly better fidelity-verifiability Pareto frontier, i.e., for any required fidelity (STS score), Waterfall (Ours) has higher verifiability than Waterfall (KGW).
@@
Another type of white-box LLM watermarking introduced in Wang et al. (2023b) modifies the model structure, and fine-tune the LLM to generate invisible Unicode watermark tokens during the generation process. This setting is extremely restrictive, as only the particular fine-tuned watermark model trained to generate the watermark tokens can be used. Additionally, similar to other Unicode watermarking methods, the watermark can be easily erased, rendering the method ineffective.
@@
L.6 Additional implementation methods like Beam search
Instead of performing conventional sampling of the perturbed logits during the watermarking text generation phase, we could apply beam search for generation. This allows us to generate several alternative versions of the watermarked text, while ensuring that the tokens being selected retains fidelity and achieves verifiability by selecting high scoring logits with strong watermark perturbations. This provides a more computationally efficient alternative to the multi-round refinement in Section L.5. Finally, the final watermarked text can be selected from the list of generations based on some criterion that balances fidelity and verifiability. We have implemented a version of this in our code release, available at https://github.com/aoi3142/Waterfall.